# Taxonomy of periodic orbits and gravitational waves in a deformed Schwarzschild black hole spacetime

### v1

@@ Abstract
In this paper, we investigate periodic orbits of test particles around a deformed Schwarzschild black hole and the resulting gravitational waves. Firstly, we examine the properties of circular orbits and find that circular orbits could disappear when the deformation is large enough. Then, using an orbital taxonomy, we characterize various periodic orbits with a set of triples, which describes the zoom–whirl behaviours. We also calculate the gravitational waveform signals generated by different periodic orbits, revealing the influence of the deformation on the gravitational wave, which can be potentially picked up by future space-based detectors.
@@ 1 Introduction
Since the successful detection of gravitational wave signal GW150914 in 2016 by LIGO and Virgo, gravitational waves have been a new messenger of cosmic explorationLIGOScientific:2016emj; LIGOScientific:2016aoc; LIGOScientific:2016vbw; LIGOScientific:2016vlm. These oscillations of spacetime were predicted by Einstein’s theory of general relativity and have significantly broadened the research horizon on the physics frontierBarack:2018yly; Berti:2015itd. Among all the sources for space-based gravitational wave detectors——such as TaijiHu:2017mde, TianqinTianQin:2015yph; Gong:2021gvw, and LISASchutz:1999xj; amaro2017laser; Gair:2004iv——the Extreme Mass Ratio Inspiral system (EMRI) is a crucial class of gravitational wave sources in the mHz frequency band, in which a stellar-mass object inspirals around a supermassive black holeHughes:2000ssa.
@@
To characterize EMRI orbits, the study of particle trajectories beyond circular orbits around black holes is of great significanceCardoso:2020iji. It helps to understand the dynamics around black holes, providing a powerful theoretical framework for addressing challenges in astrodynamics. Among all the particle trajectories, periodic orbits are particularly important, as all eccentric orbits will undergo approximately periodic orbits during their inspiral towards the central black holeLevin:2008mq; Levin:2009sk. Periodic orbits exhibit topological characteristics of so-called ”multi-leaf clovers” which means whirling around the central black hole and zooming out in quasi elliptical orbits Barack:2003fp; Glampedakis:2002ya. Three integers index orbits topologically. They represent scaling, rotation and vertex behavior, respectivelyLevin:2008mq. Therefore, an orbital taxonomy can associate periodic orbits with rational numbers. Using this kind of classification, a series of discussions have already been carried out in different spacetimes, such as Schwarzschild and KerrLevin:2008mq; Levin:2009sk; Compere:2021bkk, Reissner-NordströmPhysRevD.82.083001, quantum-corrected black holeGuo:2025scs; Yang:2024lmj, Einstein-Æther black holeLu:2025cxx, and so onLiu:2018vea; Wei:2019zdf; Gao:2020wjz; Lin:2022wda; Wang:2022tfo; Lin:2023rmo; Tu:2023xab; Jiang:2024cpe; Zahra:2025tdo; Deng:2025wzz; Lin:2021noq; Deng:2020hxw; Melis:2025iaw; Gong:2025mne; Ahmed:2025azu; Sharipov:2025yfw; Alloqulov:2025ucf. An effective way to display these orbits is to calculate gravitational waveforms, and in some cases waveforms have been studied before, see Refs.Lu:2025cxx; Yang:2024lmj; Zahra:2025tdo; Deng:2025wzz; Gong:2025mne; Ahmed:2025azu and references therein.
@@
Periodic orbits not only play an important role in orbit taxonomy but also have significance in studying transient resonances. As mentioned before, generic bound orbits can be approximately described by periodic orbits with perturbationLevin:2008mq. In Kerr spacetime, when radiation reaction is considered, small compact objects will go through low order periodic orbits and transient resonances as a post adiabatic effect will emerge, accompanied by mutations of the motion constants Flanagan:2010cd; Berry:2016bit. Moreover, resonances are focused more often in non-Kerr spacetime, where the Carter symmetry is broken and the spacetime becomes non-integrable. In this case, the resonance islands will significantly generate within the phase spacePan:2023wau.
@@
Among various studies on non-Kerr spacetime, a metric, developed by Destounis, Suvorov and KokkotasPhysRevD.102.064041 (DSK), includes generic deviations from the Kerr spacetime with a parameter $\alpha_{Q}$ controlling the Carter symmetry. In their work, gravitational-wave glitches associated with $r$-$\theta$ resonant orbits have been studiedDestounis:2021mqv; PhysRevD.102.064041. However, a systematic classification of periodic orbits in this spacetime remains to be studied. Given the complexity of non-Kerr spacetime, comprehensive analysis of periodic orbits in the generic case of this kind of spacetime is a big challenge. Therefore, we explore the $r$-$\varphi$ periodic orbits of this spacetime in static limit, namely a deformed Schwarzschild spacetime. Under this circumstance, the null circular orbit can be analytically solved. Actually, we solve null and time-like circular orbits in this spacetime and find that the circular orbits will be qualitatively different from those in the Schwarzschild spacetime in some situation. We also give a relatively complete taxonomy of periodic orbits with a triplet $(z,\omega,v)$ describing the number of orbital branches, the orbits whirling and the movement of the small object between apastrons, respectively. Based on this taxonomy, we study how the deformation affects the shape of them. At last, we also calculate the gravitational waveform signals generated by different periodic orbits, analyze the effects of deformation on them and show the differences from Schwarzschild spacetime by calculating the mismatch, which can be used to test the DSK model in future space-based observation.
@@ 2 The deformed Schwarzschild spacetime
with a parameter $\alpha$ describing the degree of deformation, which is the parameter $\alpha_{Q}$ in the non-Kerr spacetimePhysRevD.102.064041. Note that the event horizon is fixed at $r=2M$, independent of $\alpha$. This spacetime can be taken as a form of an exact, vacuum solution of mixed scalar-$f(R)$ theories Suvorov2021.
The intuition about the physical essence of the deformation can be illustrated by comparison with existing solutions. The $tt$-component is the limit of a particular non-vacuum black hole metric introduced by Bardeen, of which the $tt$-component reads $g_{tt}=-1+2M/r-3Mq^{2}/r^{3}+\mathcal{O}(1/r^{4})$. The Bardeen solution involves a metric coupled to a nonlinear electromagnetic field which is generated by Bardeen magnetic monopole. Therefore, the parameter $\alpha$ can be explained as a gravitational analogy of a Bardeen magnetic monopolePhysRevD.102.064041. More detailed analysis needs tools like multipole moment expansionsThorne:1980ru.
@@ 4 Periodic Orbits
In this section, a detailed discussion on the performance of periodic orbits will be conducted. A periodic orbit means that this trajectory can return to its initial state within a limited time. In the determination of the trajectories, this kind of orbits is very crucial. Generic orbit can be approximated by nearby periodic orbits because any irrational number can be infinitely approximated by rational numbers. Therefore, so it can help to analyze the evolution of EMRIs.
@@ 5 Gravitational Radiation from periodic orbits
The numerical Kludge model will be introduced briefly primarily, since it captures many qualitative features of EMRIsChua:2017ujo; Babak:2006uv. It can even be available in self-force condition by including the particle dynamics. Although the self-consistency will degrade, it is still effective as an approximation to Teukolsky-based waveform in strong fieldChua:2017ujo. To compute the waveform, equations of motion (4) need be solved numerically to construct trajectory in space of motion constants firstly. As mentioned before, the influence caused by radiation reaction can be ignored. Secondly, the curved-space coordinates of trajectory then will be associated artificially with coordinates in flat space. Using the standard quadrupole formula, the corresponding gravitational waveform is generated.
The symmetric trace-free (STF) mass quadrupole of the small object with mass $m$ isThorne:1980ru:
@@
where $\iota$ and $\eta$ represent the inclination angle and the longitude of the perihelion of the orbital plane, respectivelyBarack:2003fp. Then the gravitational wave polarizations can be obtained as follows:
@@
Considering the transverse-traceless gauge, the concrete expressions of crucial components are given asBabak:2006uv:
@@
In order to show the influence of the deformation parameter on the gravitational waveforms of different periodic orbits, we here consider a small compact object with mass $m=10M_{\odot}$ around a supermassive black hole with mass $M=10^{6}M_{\odot}$ at the luminosity distance $D_{\text{L}}=1\text{Gpc}$. Take a simple case, the inclination angle $\iota$ is set as $\pi/2$ and the longitude of perihelion is fixed at 0. Moreover, the polarization angle is also set as 0 for simplicityBabak:2006uv. To show contrast, we fix the energy at $E=\sqrt{0.9755751102284231}$ and change the deviation parameter $\alpha$. When $\alpha$ is large enough, the perihelion is too small, as shown in Fig.6. So numerical Kludge model will not work. And then the waveforms are represented by two polarization components $h_{+}$ and $h_{\times}$ to show their characteristics in Fig.7. We list three parameter conditions and their waveform contrast here.
@@
In Figs6 and 7, a typical kind of periodic orbits is plotted with the triple $(1,2,0)$ and their $h_{+}$ and $h_{\times}$ modes with different deformation parameters. It shows that a clear relationship exists between gravitational waveforms and orbit. The giant leave results in very long smooth phases. And there are rapid oscillations which align with the zoom and whirl behaviors. Moreover, it is clear that in comparison with the Schwarzschild case the parameter influences the phase of the waveforms mainly and the waveform amplitudes in a slight way when parameter increases. To clearly illustrate the differences in gravitational waveforms with different parameters, a useful characteristic is the mismatch between to waveformsAllen:2019vcl; Qiao:2024gfb; Lai:2025zfm. To conduct a preliminary analysis of the differences, we calculated the mismatch according to the Schwarzschild condition with $\alpha$ from 0 to 0.4 and concluded the trend below in Fig.8:
@@
We find that the mismatch will be over $0.1$ when $\alpha\geq 0.5$, so we don’t consider it. The value of mismatch, apparently, goes up with the increase of the deformation parameter. It is reasonable because it means that difference from the Schwarzschild situation increases. We can preliminarily distinguish the difference by the value of mismatch accumulating over time although the effect of the deformation is relatively weak in the amplitude and phases. Certainly, further detailed differentiation of gravitational waveforms needs more detailed detectability and measurability analysis. Further research can be conducted based on the reference to Speeney:2022ryg. These findings have pointed the way of detecting detailed physical characteristics in the future.
@@ 6 Conclusions and Discussions
The work presented here reveals some distinct features of geodesics from the Schwarzschild black hole. Nevertheless, there are some intriguing problems to address in the future. For example, orbital resonances can be taken into further investigationFlanagan:2010cd. Moreover, the possible mechanism in the inner region and the disappearance of the photon ring are also worth considering. Last but not the least, while research for geodesics in DSK metric under general conditions is challenging, qualitative analysis can be taken into consideration. We leave these topics to future work.

### v2

@@ Abstract
In this paper, we investigate periodic orbits of test particles around a deformed Schwarzschild black hole and the resulting gravitational waves. Firstly, we examine the properties of circular orbits and find that circular orbits could disappear when the deformation is large enough. Then, using an orbital taxonomy, we characterize various periodic orbits with a set of triples, which describes the zoom-whirl behaviours. We also calculate the gravitational waveform signals generated by different periodic orbits, revealing the influence of the deformation on the gravitational wave, which can be potentially picked up by future space-based detectors.
@@ 1 Introduction
Since the successful detection of gravitational wave signal GW150914 in 2016 by LIGO and Virgo, gravitational waves have been a new messenger of cosmic explorationAbbott and others (2016b, c, a, d). These oscillations of spacetime were predicted by Einstein’s theory of general relativity and have significantly broadened the research horizon on the physics frontierBarack and others (2019); Berti and others (2015). Among all the sources for space-based gravitational wave detectors——such as TaijiHu and Wu (2017), TianqinLuo and others (2016); Gong et al. (2021), and LISASchutz (1999); Amaro-Seoane et al. (2017); Gair et al. (2004)——the Extreme Mass Ratio Inspiral system (EMRI) is a crucial class of gravitational wave sources in the mHz frequency band, in which a stellar-mass object inspirals around a supermassive black holeHughes (2001).
@@
To characterize EMRI orbits, the study of particle trajectories beyond circular orbits around black holes is of great significanceCardoso et al. (2021). It helps to understand the dynamics around black holes, providing a powerful theoretical framework for addressing challenges in astrodynamics. Among all the particle trajectories, periodic orbits are particularly important, as all eccentric orbits will undergo approximately periodic orbits during their inspiral towards the central black holeLevin and Perez-Giz (2008); Levin (2009). Periodic orbits exhibit topological characteristics of so-called ”multi-leaf clovers” which means whirling around the central black hole and zooming out in quasi elliptical orbits Barack and Cutler (2004); Glampedakis and Kennefick (2002). Three integers index orbits topologically. They represent scaling, rotation and vertex behavior, respectivelyLevin and Perez-Giz (2008). Therefore, an orbital taxonomy can associate periodic orbits with rational numbers. Using this kind of classification, a series of discussions have already been carried out in different spacetimes, such as Schwarzschild and KerrLevin and Perez-Giz (2008); Levin (2009); Compère et al. (2022), Reissner-NordströmMisra and Levin (2010), quantum-corrected black holeGuo et al. (2025); Yang et al. (2025), Einstein-Æther black holeLu and Zhu (2025), and so onLiu et al. (2019); Wei et al. (2019); Gao and Deng (2020); Lin and Deng (2022); Wang et al. (2022); Lin and Deng (2023); Tu et al. (2023); Jiang et al. (2024); Zahra et al. (2025); Deng et al. (2025); Lin and Deng (2021); Deng (2020); Melis et al. (2025); Gong et al. (2025); Ahmed et al. (2025a); Sharipov et al. (2025); Alloqulov et al. (2025). An effective way to display these orbits is to calculate gravitational waveforms, and in some cases waveforms have been studied before, see Refs.Lu and Zhu (2025); Yang et al. (2025); Zahra et al. (2025); Deng et al. (2025); Gong et al. (2025); Ahmed et al. (2025a, b) and references therein.
@@
Periodic orbits not only play an important role in orbit taxonomy but also have significance in studying transient resonances. As mentioned before, generic bound orbits can be approximately described by periodic orbits with perturbationLevin and Perez-Giz (2008). In Kerr spacetime, when radiation reaction is considered, small compact objects will go through low order periodic orbits and transient resonances as a post adiabatic effect will emerge, accompanied by mutations of the motion constants Flanagan and Hinderer (2012); Berry et al. (2016). Moreover, resonances are focused more often in non-Kerr spacetime, where the Carter symmetry is broken and the spacetime becomes non-integrable. In this case, the resonance islands will significantly generate within the phase spacePan et al. (2023).
@@
Among various studies on non-Kerr spacetime, a metric, developed by Destounis, Suvorov and KokkotasDestounis et al. (2020) (DSK), includes generic deviations from the Kerr spacetime with a parameter $\alpha_{Q}$ controlling the Carter symmetry. In their work, gravitational-wave glitches associated with $r$-$\theta$ resonant orbits have been studiedDestounis et al. (2021, 2020). However, a systematic classification of periodic orbits in this spacetime remains to be studied. Given the complexity of non-Kerr spacetime, comprehensive analysis of periodic orbits in the generic case of this kind of spacetime is a big challenge. Therefore, we explore the $r$-$\varphi$ periodic orbits of this spacetime in static limit, namely a deformed Schwarzschild spacetime. Under this circumstance, the null circular orbit can be analytically solved. Actually, we solve null and time-like circular orbits in this spacetime and find that the circular orbits will be qualitatively different from those in the Schwarzschild spacetime in some situation. We also give a relatively complete taxonomy of periodic orbits with a triplet $(z,w,v)$ describing the number of orbital branches, the orbits whirling and the movement of the small object between apastrons, respectively. Based on this taxonomy, we study how the deformation affects the shape of them. At last, we also calculate the gravitational waveform signals generated by different periodic orbits, analyze the effects of deformation on them and show the differences from Schwarzschild spacetime by calculating the mismatch, which can be used to test the DSK model in future space-based observation.
@@ 2 The deformed Schwarzschild spacetime
with a parameter $\alpha$ describing the degree of deformation, which is the parameter $\alpha_{Q}$ in the non-Kerr spacetimeDestounis et al. (2020). Note that the event horizon is fixed at $r=2M$, independent of $\alpha$. This spacetime can be taken as a form of an exact, vacuum solution of mixed scalar-$f(R)$ theories Suvorov (2021).
The intuition about the physical essence of the deformation can be illustrated by comparison with existing solutions. The $tt$-component is the limit of a particular non-vacuum black hole metric introduced by Bardeen, of which the $tt$-component reads $g_{tt}=-1+2M/r-3Mq^{2}/r^{3}+\mathcal{O}(1/r^{4})$. The Bardeen solution involves a metric coupled to a nonlinear electromagnetic field which is generated by Bardeen magnetic monopole. Therefore, the parameter $\alpha$ can be explained as a gravitational analogy of a Bardeen magnetic monopoleDestounis et al. (2020). More detailed analysis needs tools like multipole moment expansionsThorne (1980).
@@ 4 Periodic Orbits
In this section, a detailed discussion on the performance of periodic orbits will be conducted. A periodic orbit means that this trajectory can return to its initial state within a limited time. In the determination of the trajectories, this kind of orbits is very crucial. Generic orbit can be approximated by nearby periodic orbits because any irrational number can be infinitely approximated by rational numbers. Therefore, it can help to analyze the evolution of EMRIs.
@@ 5 Gravitational Radiation from periodic orbits
The numerical Kludge model will be introduced briefly primarily, since it captures many qualitative features of EMRIsChua et al. (2017); Babak et al. (2007). It can even be available in self-force condition by including the particle dynamics. Although the self-consistency will degrade, it is still effective as an approximation to Teukolsky-based waveform in strong fieldChua et al. (2017). To compute the waveform, equations of motion (4) need be solved numerically to construct trajectory in space of motion constants firstly. As mentioned before, the influence caused by radiation reaction can be ignored. Secondly, the curved-space coordinates of trajectory then will be associated artificially with coordinates in flat space. Using the standard quadrupole formula, the corresponding gravitational waveform is generated.
The symmetric trace-free (STF) mass quadrupole of the small object with mass $m$ isThorne (1980):
@@
where $\iota$ and $\eta$ represent the inclination angle and the longitude of the perihelion of the orbital plane, respectivelyBarack and Cutler (2004). Then the gravitational wave polarizations can be obtained as follows:
@@
Considering the transverse-traceless gauge, the concrete expressions of crucial components are given asBabak et al. (2007):
@@
In order to show the influence of the deformation parameter on the gravitational waveforms of different periodic orbits, we here consider a small compact object with mass $m=10M_{\odot}$ around a supermassive black hole with mass $M=10^{6}M_{\odot}$ at the luminosity distance $D_{\text{L}}=1\text{Gpc}$. Take a simple case, the inclination angle $\iota$ is set as $\pi/2$ and the longitude of perihelion is fixed at 0. Moreover, the polarization angle is also set as 0 for simplicityBabak et al. (2007). To show contrast, we fix the energy at $E=\sqrt{0.9755751102284231}$ and change the deviation parameter $\alpha$. When $\alpha$ is large enough, the perihelion is too small, as shown in Fig.6. So numerical Kludge model will not work. And then the waveforms are represented by two polarization components $h_{+}$ and $h_{\times}$ to show their characteristics in Fig.7. We list three parameter conditions and their waveform contrast here.
@@
In Figs6 and 7, a typical kind of periodic orbits is plotted with the triple $(1,2,0)$ and their $h_{+}$ and $h_{\times}$ modes with different deformation parameters. It shows that a clear relationship exists between gravitational waveforms and orbit. The giant leave results in very long smooth phases. And there are rapid oscillations which align with the zoom and whirl behaviors. Moreover, it is clear that in comparison with the Schwarzschild case the parameter influences the phase of the waveforms mainly and the waveform amplitudes in a slight way when parameter increases. To clearly illustrate the differences in gravitational waveforms with different parameters, a useful characteristic is the mismatch between to waveformsAllen (2019); Qiao et al. (2025); Lai et al. (2025). To conduct a preliminary analysis of the differences, we calculated the mismatch according to the Schwarzschild condition with $\alpha$ from 0 to 0.4 and concluded the trend below in Fig.8:
@@
We find that the mismatch will be over $0.1$ when $\alpha\geq 0.5$, so we don’t consider it. The value of mismatch, apparently, goes up with the increase of the deformation parameter. It is reasonable because it means that difference from the Schwarzschild situation increases. We can preliminarily distinguish the difference by the value of mismatch accumulating over time although the effect of the deformation is relatively weak in the amplitude and phases. Certainly, further detailed differentiation of gravitational waveforms needs more detailed detectability and measurability analysis. Further research can be conducted based on the reference to Speeney et al. (2022). These findings have pointed the way of detecting detailed physical characteristics in the future.
@@ 6 Conclusions and Discussions
The work presented here reveals some distinct features of geodesics from the Schwarzschild black hole. Nevertheless, there are some intriguing problems to address in the future. For example, orbital resonances can be taken into further investigationFlanagan and Hinderer (2012). Moreover, the possible mechanism in the inner region and the disappearance of the photon ring are also worth considering. Last but not the least, while research for geodesics in DSK metric under general conditions is challenging, qualitative analysis can be taken into consideration. We leave these topics to future work.